How many liters of a 15% acid solution must be mixed with 5 liters of a 25% acid solution to give a 20% solution?
Answer by stanbon(75887)

How many liters of a 15% acid solution must be mixed with 5 liters of a 25% acid solution to give a 20% solution?
----
Equation:
acid + acid = acid
0.15x + 0.25*5 = 0.20(x+5)
--------
15x + 25*5 = 20*x + 20*5
--------
-5x = -5*5
----
x = 5 liters (amt of 15% solution needed)
=====================
Cheers,
Stan H.
-----------------------